# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ VII.4 DeepH
The deep-learning density functional theory Hamiltonian (DeepH) method is a neural network approach based on equivariant graph neural networks for modeling the DFT Hamiltonian as a function of material structure Li et al. (2022b). Leveraging the sparsity of the Hamiltonian matrix under NAOs and its compatibility with Walter Kohn’s “quantum nearsightness principle" Kohn (1996), the DeepH method can learn from training data of small structures to infer electronic Hamiltonians of large structures, achieving high prediction accuracy with linear-scaling computational cost. The predicted Hamiltonian may be subsequently utilized for post-processing to evaluate properties including band structures, optical properties, and response properties from density functional perturbation theory, etc Li et al. (2022b); Gong et al. (2023); Li et al. (2024b).
@@
The interface between ABACUS and DeepH was created in 2022, shortly after the invention of the DeepH approach in 2021. The interface ensures compatibility between ABACUS and various versions of DeepH, including DeepH-pack Li et al. (2022b); DeepH , DeepH-E3 Gong et al. (2023); DeepH-E (3), and DeepH-2 Wang et al. (2024a). Among these, DeepH-E3 stands out as the most stable open-source implementation to date and has been utilized with ABACUS in several example studies DeepH-E (3); Tang et al. (2024); Yang et al. (2024). The ABACUS-DeepH interface is versatile, supporting spin-orbit coupling and magnetic systems, and facilitates ABACUS’s integration with xDeepH–a specialized variant of DeepH for predicting the electronic structures of magnetic materials Li et al. (2023). To address the modified sparsity patterns in hybrid DFT Hamiltonians, a dedicated toolkit, named DeepH-hybrid, has been developed, which verified DeepH’s applicability to hybrid-functional Hamiltonians generated by ABACUS Tang et al. (2024, 2023). An interface between ABACUS and DeepH-DFPT (a generalization of the DeepH approach for deep-learning density functional perturbation theory) is currently under development to accelerate calculations of electron-phonon coupling Li et al. (2024b). In addition, a recently developed “Hamiltonian Projection and Reconstruction to atomic Orbitals" (H-PRO) method can transform the DFT Hamiltonian of plane-wave basis into localized basis, making the DeepH compatible with the plane-wave mode of ABACUS Gong et al. (2024). Very recently, a universal materials model (UMM) of DeepH, named DeepH-UMM, has been developed, demonstrating exceptional transferability across a wide range of material structures composed of various elements Wang et al. (2024b). Given ABACUS’s versatility, further development of DeepH-UMM in conjunction with ABACUS holds significant promise for advancing materials discovery.
@@
The current ABACUS-DeepH interface is available on GitHub DeepH . For data preparation, the keyword out_mat_hs2=1 is mandatory in the ABACUS input file, enabling the formatted output of the Hamiltonian and overlap matrix in a sparse matrix (CSR) form. For inference with trained DeepH models, ABACUS’s get_S function provides an efficient method for generating the overlap matrix, from which the Hamiltonian can be predicted by DeepH. It is important to note the overlap and Hamiltonian matrices are assumed to shared the same sparsity pattern in DeepH. To ensure clarity, fixing such sparsity pattern with tools provided in DeepH-hybrid is recommended Tang et al. (2023). Tutorials for the interface are available online, and four demo datasets with hybrid DFT Hamiltonians computed with ABACUS are publicly available on Zenodo DeepH ; Tang et al. (2023).
VII.5 DeePTB
@@
DeePTB is an innovative open-source package that leverages deep learning to accelerate ab initio electronic simulationsGu et al. (2024); Zhouyin et al. (2024). Its integration with ABACUS creates a powerful synergy, combining first-principles calculations with advanced machine learning techniques for efficient large-scale electronic structure predictions. Notably, it has demonstrated the capability to simulate systems containing up to millions of atomsGu et al. (2024), a scale previously unattainable with traditional methods.
@@
Fig. 38 illustrates the ABACUS-DeePTB workflow. ABACUS performs DFT calculations on training structural data obtained from crystal databases or molecular dynamics simulations, generating essential quantum mechanical data as labels including energy eigenvalues, NAOs basis Hamiltonian, overlap, and density matrices. DeePTB utilizes this ABACUS-generated data to train two key models: the DeePTB-SK Gu et al. (2024) and DeePTB-E3 Zhouyin et al. (2024) models. The DeePTB-SK model uses energy eigenvalues to develop an improved Slater-Koster TB models for efficient electronic structure prediction, while the DeePTB-E3 model, based on the SLEM (Strictly Localized Equivariant Message-passing) approach, predicts quantum operators including the Hamiltonian, overlap, and density matrices.
@@ VII.6 PyATB
PYATB (Python ab initio tight-binding simulation package) is a Python package based on ab initio tight-binding Hamiltonian, designed as a tool for calculating and analyzing the electronic band structures of materialsJin et al. (2023). It can be viewed as a post-processing program for the ABACUS. When ABACUS completes the self-consistent electronic calculations and generates the tight-binding Hamiltonian, PYATB utilizes this Hamiltonian to perform electronic property calculations. This eliminates the need for the cumbersome construction of MLWFs while strictly preserving the Hamiltonian’s symmetry, making it particularly well-suited for high-throughput workflows in electronic structure calculations of various materials. Currently, PYATB mainly offers three functional modules: Bands, Geometric, and Optical. This allows ABACUS, in combination with PYATB, to compute basic band structures, fat bands, and projected density of states (PDOS). Additionally, it can analyze the Berry curvature and Chern number of topological materialsXiao, Chang, and Niu (2010); Hasan and Kane (2010), as well as compute first- and second-order optical responses such as optical conductivity, shift current, second harmonic generationSturman and Fridkin (1992); Sipe and Shkrebtii (2000); Jin and He and Berry curvature dipoleSodemann and Fu (2015); Pang, Jin, and He .
ABACUS standard self-consistent calculations can generate the Hamiltonian matrix $\mathbf{H}(\mathbf{k})$, the overlap matrix $\mathbf{S}(\mathbf{k})$, and the dipole matrix
$\mathbf{A}_{\mu\nu}^{\mathbf{R}}(\mathbf{k})=\sum_{\mathbf{R}}\mathrm{e}^{i\mathbf{k}\cdot\mathbf{R}}\langle\phi_{\mu 0}|\mathbf{r}|\phi_{\nu\mathbf{R}}\rangle.$ (141)
Using these input parameters, PYATB can solve the following generalized eigenvalue problem:
$\mathbf{H}(\mathbf{k})\mathbf{C}_{n}(\mathbf{k})=E_{n\mathbf{k}}\mathbf{S}(k)\mathbf{C}(\mathbf{k}),$ (142)
where $\mathbf{C}_{n}(\mathbf{k})$ is the eigenvector of the $n$-th band, then obtain the band structure and Bloch wavefunctions information. Subsequently, based on the wavefunctions, the Berry curvatureXiao, Chang, and Niu (2010); Jin, Zheng, and He (2021) can be calculated as:
$\Omega_{n}(\mathbf{k})=\nabla_{\mathbf{k}}\times i\langle u_{n\mathbf{k}}|\nabla_{\mathbf{k}}|u_{n\mathbf{k}}\rangle.$ (143)
This allows for the calculation of various topological and optical properties of materials. One can refer to Ref. Jin et al., 2023 for the implementation details.
VII.7 Hefei-NAMD
@@
Hefei Non-Adiabatic Molecular Dynamics (Hefei-NAMD) is an ab initio simulation suite for studying excited carrier dynamics in condensed matter systems. It combines real-time time-dependent density functional theory (TDDFT) with the fewest-switches surface hopping scheme and classical-path approximation. Hefei-NAMD has been used to investigate processes such as charge transfer, electron–hole recombination, spin dynamics and exciton dynamicsZheng et al. (2019); Zheng, Zheng, and Zhao (2022); Jiang et al. (2021); Zheng et al. (2023). It allows the study of excited carrier dynamics in energy, real, and momentum spaces, while also exploring interactions with phonons, defects, and molecular adsorptions, providing valuable insights into ultrafast dynamics at the atomic scale. Hefei-NAMD works by interfacing with other ab initio codes, e.g. VASP, ABACUS etc. Figure 39 shows the flowchart of the simulation using Hefei-NAMD interfaced with ABACUS. Detail documentations can be found on the ABACUS webciteaba (2024c).
@@
VII.8 PEXSI
PEXSI (Pole EXpansion and Selected Inversion) Lin et al. (2009, 2013) provides an alternative method for solving the KS equation without using diagonalization methods. Given the generalized eigenvalue problem $H\Phi=\varepsilon S\Phi$ transferred from KS equation, PEXSI directly calculates the real-space density matrix from $H$ and $S$ matrix.
We denote the collection of all atomic orbitals in the real space $\Phi\left(\mathbf{r}\right)$ as $\left[\phi_{1}(\mathbf{r}),\cdots,\phi_{M}(\mathbf{r})\right]$, and the single particle real-space density matrix $\hat{\gamma}\left(\mathbf{r},\mathbf{r^{\prime}}\right)$ can be formally expressed as
$\hat{\gamma}\left(\mathbf{r},\mathbf{r^{\prime}}\right)=\Phi\left(\mathbf{r}\right)f(H-\mu S)\Phi^{*}\left(\mathbf{r^{\prime}}\right),$ (144)
where $f$ is the Fermi-Dirac distribution function and $\mu$ is the chemical potential. From the density matrix, we can get charge density in the following form:
$\rho\left(\mathbf{r}\right)=\hat{\gamma}\left(\mathbf{r},\mathbf{r}\right).$ (145)
PEXSI expands the matrix Fermi-Dirac function using a $P$-term pole expansion:
$f(H-\mu S)\approx\mathrm{Im}\sum_{l=1}^{P}\omega_{l}^{\rho}\left[H-(z_{l}+\mu)S\right]^{-1},$ (146)
where coefficients $z_{l}$ and $\omega_{l}$ are chosen carefully and can be computed in advance. PEXSI uses selected inversion to calculate these inverse matrices effectively.
@@ VII.8 PEXSI
The computational cost of standard diagonalization methods scales as $\mathcal{O}(N^{3})$, while PEXSI scales at most as $\mathcal{O}(N^{2})$, where $N$ is the dimension of the matrix. The PEXSI method has a two-level parallelism structure (pole expansion parallelism and selected inversion parallelism) and is designed to be of high scalability. The sequential version of PEXSI has been tested before with ABACUSLin et al. (2013). Now, ABACUS has implemented an interface to the parallel version of PEXSI for solving the KS equation on NAO basis.
VII.9 Others
The Atomic Simulation Environment (ASE) comprises a comprehensive suite of tools and Python modules designed for setting up, manipulating, executing, visualizing, and analyzing atomic simulationsASE (2024a). In our framework, we have developed an ABACUS calculatorASE (2024b) that interfaces with the latest version of ASE (3.23.0b1) and that enables seamless integration of ASE’s robust functionalities for both pre-processing and post-processing tasks.
Regarding pre-processing capabilities, the interface facilitates tasks such as converting structural files a in various formats, generating uniform k-point grids tailored to specific simulation needs, and establishing high-symmetry paths essential for accurate calculations of electronic band structures and other properties. For post-processing, the integration with ASE empowers ABACUS users with advanced tools for visualizing and analyzing simulation results. This includes visualizing the structure and energy plot of the optimization process, plotting electronic band structures, and performing MD trajectory analyses.
@@ VII.9 Others
Moreover, the collaboration between ABACUS and ASE extends beyond basic simulations to encompass complex computational tasks. The self-consistent calculations of ABACUS are combined with ASE’s built-in optimization algorithms, facilitating structural relaxation procedures, conducting MD simulations under varying conditions, performing precise phonon calculations based on the finite displacement method, carrying-out global optimization by genetic algorithm (GA), locating transition states (TS) of reactions by saddle point refinement, and more.
For TS search, as an example, two most widely used methods, the nudged elastic band (NEB)Jónsson, Mills, and Jacobsen (1998); Henkelman and Jónsson (2000); Henkelman, Uberuaga, and Jónsson (2000) method and the DimerHenkelman and Jónsson (1999); Olsen et al. (2004); Heyden, Bell, and Keil (2005); Kästner and Sherwood (2008) method, which represent double-ended TS searching methods and single-end TS methods respectively, are implemented in ASE. Besides, some other enhanced NEB methods are coded in the ASE package, such as dynamic NEB Lindgren, Kastlunger, and Peterson (2019) and AutoNEB Kolsbjerg, Groves, and Hammer (2016). Furthermore, there are many other saddle point refinement algorithms implemented in ASE, and some of them show better performance, like Sella Hermes et al. (2019, 2021, 2022). With the above-mentioned methods available in ASE, ABACUS can be used for searching for transition states.
Apart from directly using TS search tools in ASE, the cooperation of different TS search methods can be done due to ASE’s coding flexibility, which leads to better TS locating functionality. One way is to first generate a rough TS by NEB, then utilize a single-ended method like Dimer or Sella to optimize the TS in the target threshold, which incorporates the advantage of both methods. As is shown in Fig. 40, ATST-Tools scripts suite ATS (2024) is prepared for handy usage of ABACUS and ASE in TS locating jobs.
@@
Phonopy Togo et al. (2023); Togo (2023) is a versatile open-source software package designed for calculating phonon and related lattice dynamics properties from first-principles calculations. With the addition of the ABACUS interface in Phonopy version 2.19.1, users can now leverage the computational capabilities of ABACUS to compute the electronic structure and force constants that Phonopy requires for its phonon calculations.
The phonon dispersion relation describes the relationship between the energy of phonons (lattice vibrations) and their wave vector ($\mathbf{q}$) within a crystal. Under the harmonic approximation, the energy of a phonon mode at a given point, $\mathbf{q}$, and the polarization index $\nu$ can be calculated as follows:
$\hbar\omega_{\mathbf{q}\nu}=\sqrt{\frac{\hbar^{2}}{2\mu}\sum_{i,j}M_{i}M_{j}|\mathbf{e}_{\mathbf{q}\nu,i}^{*}\cdot{D}_{ij}(\mathbf{q})\cdot\mathbf{e}_{\mathbf{q}\nu,j}|^{2}},$ (147)
where $\omega_{\mathbf{q}\nu}$ is the angular frequency of the phonon mode; $D_{ij}$$(\mathbf{q})$ is the dynamical matrix element between atoms $i$ and $j$ at the $\mathbf{q}$-point; $\mathbf{e}_{\mathbf{q}\nu,i}$ is the eigenvector associated with atom $i$ for the $\nu$th phonon mode at $\mathbf{q}$; $M_{i}$ and $M_{j}$ are the masses of atoms $i$ and $j$; $\mu$ is the reduced mass, typically $M_{i}M_{j}/(M_{i}+M_{j})$.
${D}_{ij}(\mathbf{q})$ is derived from the force constants, which relate the forces on each atom to the displacements of other atoms in the crystal. ABACUS can provide the force constants through finite displacement methods, where small displacements are applied to atoms in the crystal. Recent research efforts Wang et al. (2020); Liu et al. (2024b) highlighted the success on the synergy between ABACUS and Phonopy to explore novel materials and optimize existing ones utilizing accurate phonon spectra.
@@
TB2JHe et al. (2021) is an open-source Python package for automatic computation of magnetic interactions between atoms of magnetic crystals from density functional Hamiltonians based on Wannier functions or linear combinations of atomic orbitals. The program is based on Green’s function method with the local rigid spin rotation treated as a perturbationLiechtenstein et al. (1987). The ABACUS interface has been added since TB2J version 0.8.0. With the Hamiltonian matrix $\mathbf{H}(\mathbf{k})$ and the overlap matrix $\mathbf{S}(\mathbf{k})$ obtained by an LCAO calculation, TB2J is able to compute the parameters of the Heisenberg Hamiltonian
$H_{\mathrm{H}}=-\sum_{i\neq j}J_{ij}^{iso}\overrightarrow{S}_{i}\cdot\overrightarrow{S}_{j}-\sum_{i\neq j}\overrightarrow{S}_{i}\cdot\mathbf{J}_{ij}^{ani}\cdot\overrightarrow{S}_{j}-\sum_{i\neq j}\overrightarrow{D}_{ij}\cdot(\overrightarrow{S}_{i}\times\overrightarrow{S}_{j}),$ (148)
where $J_{ij}^{iso}$ represents the isotropic exchange, $\mathbf{J}_{ij}^{ani}$ represents the symmetric anisotropic exchange which is a 3$\times$3 tensor, $\overrightarrow{D}_{ij}$ represents the Dzyaloshinskii-Moriya interaction (DMI). In addition, one can add the single-ion anisotropy (SIA) into the above Hamiltonian. In 2024, Zhang et al. Zhang et al. (2024e) calculate the $J_{ij}^{iso}$, $\mathbf{J}_{ij}^{ani}$ and $\overrightarrow{D}_{ij}$ for GeFe3N by using ABACUS--TB2J, which help them study the short-range order and strong interplay between local and itinerant magnetism in GeFe3N.
@@
Wannier90Pizzi et al. (2020) is a software package designed for calculating maximally-localized Wannier functions (MLWFs)Marzari and Vanderbilt (1997), which facilitates cost-effective band structure calculations. MLWFs are often used to construct model Hamiltonians. The post-processing capabilities of Wannier90 enable the calculation of various properties, such as Berry curvature, Berry curvature dipole, and shift currentPizzi et al. (2020). Additionally, MLWFs generated by Wannier90 can be analyzed for material surface states and topological properties using the Wannier ToolsWu et al. (2018).
According to the standard scheme for generating maximally localized Wannier functions in Wannier90, ABACUS has implemented an interface compatible with the Wannier90 software. This interface allows ABACUS to generate the necessary files for Wannier90, including overlap files (*.mmn) between the periodic parts of Bloch functions at neighboring $\mathbf{k}$-points, projection files (*.amn) of Bloch functions onto trial localized orbitals, eigenvalue files (*.eig) of the energy bands, and real-space distribution files (UNK.*) of the periodic parts of Bloch functions.
In the numerical atomic orbitals basis calculation, the *.mmn files are calculated using radial Gauss–Legendre and angular Lebedev-Laikov quadrature gridsLebedev and Laikov (1999). The overlap matrix elements are computed as follows:
$\langle u_{n\mathbf{k}}|$ $u_{m,\mathbf{k}+\mathbf{b}}\rangle$ (149)
$=\sum_{\mu\nu}C_{n\mu}(\mathbf{k})^{*}C_{m\nu}(\mathbf{k}+\mathbf{b})\sum_{\mathbf{R}}\mathrm{e}^{i\left(\mathbf{k}+\mathbf{b}\right)\cdot\mathbf{R}}\langle\phi_{\mu 0}|\mathrm{e}^{-i\mathbf{b}\cdot\mathbf{r}}|\phi_{\nu\mathbf{R}}\rangle,$
where $u_{n\mathbf{k}}$ is the periodic part of the Bloch wave function $\psi_{n\mathbf{k}}$, $\mathbf{k}$ and $\mathbf{k}+\mathbf{b}$ are two neighboring points, and $C_{n\mu}$ represents the expansion coefficients of $\psi_{n\mathbf{k}}$ in terms of atomic orbitals $\phi_{\mu}$. This method ensures precise calculation of the overlap matrix elements, which are essential for constructing accurate MLWFs.
VIII Summary

### v3

@@ VI.4 DeepH
The deep-learning density functional theory Hamiltonian (DeepH) method is a neural network approach based on equivariant graph neural networks for modeling the DFT Hamiltonian as a function of material structure Li et al. (2022b). Leveraging the sparsity of the Hamiltonian matrix under NAOs and its compatibility with Walter Kohn’s “quantum nearsightness principle" Kohn (1996), the DeepH method can learn from training data of small structures to infer electronic Hamiltonians of large structures, achieving high prediction accuracy with linear-scaling computational cost. The predicted Hamiltonian may be subsequently utilized for post-processing to evaluate properties including band structures, optical properties, and response properties from density functional perturbation theory, etc Li et al. (2022b); Gong et al. (2023); Li et al. (2024).
@@
The interface between ABACUS and DeepH was developed in 2022, shortly after the invention of the DeepH approach in 2021. The interface ensures compatibility between ABACUS and various versions of DeepH, including DeepH-pack Li et al. (2022b); DeepH , DeepH-E3 Gong et al. (2023); DeepH-E (3), and DeepH-2 Wang et al. (2024a). Among these, DeepH-E3 stands out as the most stable open-source implementation to date and has been utilized with ABACUS in several example studies DeepH-E (3); Tang et al. (2024); Yang et al. (2024). The ABACUS-DeepH interface is versatile, supporting spin-orbit coupling and magnetic systems, and facilitates ABACUS’s integration with xDeepH–a specialized variant of DeepH for predicting the electronic structures of magnetic materials Li et al. (2023). To address the modified sparsity patterns in hybrid DFT Hamiltonians, a dedicated toolkit, named DeepH-hybrid, has been developed, which verified DeepH’s applicability to hybrid-functional Hamiltonians generated by ABACUS Tang et al. (2024, 2023). An interface between ABACUS and DeepH-DFPT (a generalization of the DeepH approach for deep-learning density functional perturbation theory) is currently under development to accelerate calculations of electron-phonon coupling Li et al. (2024). In addition, a recently developed “Hamiltonian Projection and Reconstruction to atomic Orbitals" (H-PRO) method can transform the DFT Hamiltonian of plane-wave basis into localized basis, making the DeepH compatible with the plane-wave mode of ABACUS Gong et al. (2024). Very recently, a universal materials model (UMM) of DeepH, named DeepH-UMM, has been developed, demonstrating exceptional transferability across a wide range of material structures composed of various elements Wang et al. (2024b). Given ABACUS’s versatility, further development of DeepH-UMM in conjunction with ABACUS holds significant promise for advancing materials discovery.
@@
The current ABACUS-DeepH interface is available on GitHub DeepH . For inference with trained DeepH models, ABACUS provides an efficient method for generating the overlap matrix, from which the Hamiltonian can be predicted by DeepH. It is important to note that the overlap and Hamiltonian matrices are assumed to share the same sparsity pattern in DeepH. To ensure clarity, fixing such sparsity pattern with tools provided in DeepH-hybrid is recommended Tang et al. (2023).
VI.5 DeePTB
@@
DeePTB is an open-source package that leverages deep learning to accelerate ab initio electronic simulations Gu et al. (2024); Zhouyin et al. (2024). Its integration with ABACUS creates a powerful synergy, combining first-principles calculations with advanced machine learning techniques for efficient large-scale electronic structure predictions. Notably, it has demonstrated the capability to simulate systems containing up to millions of atoms Gu et al. (2024), a scale previously unattainable with traditional methods.
@@
Fig. 33 illustrates the ABACUS-DeePTB workflow. ABACUS performs DFT calculations on training structural data obtained from crystal databases or molecular dynamics simulations, generating essential quantum mechanical data as labels including energy eigenvalues, NAOs basis Hamiltonian, overlap, and density matrices. DeePTB utilizes this ABACUS-generated data to train two key models: the DeePTB-SK Gu et al. (2024) and DeePTB-E3 Zhouyin et al. (2024) models. The DeePTB-SK model uses energy eigenvalues to develop an improved Slater-Koster TB model for efficient electronic structure prediction, while the DeePTB-E3 model, based on the SLEM (Strictly Localized Equivariant Message-passing) approach, predicts quantum operators including the Hamiltonian, overlap, and density matrices.
@@ VI.6 PYATB
PYATB (PYthon Ab initio Tight-Binding simulation package) is a Python package based on ab initio tight-binding Hamiltonian, designed as a tool for calculating and analyzing the electronic structures of materials Jin et al. (2023). It can be viewed as a post-processing program for ABACUS. When ABACUS completes the SCF calculations and generates the tight-binding Hamiltonian, PYATB utilizes this Hamiltonian to perform electronic property calculations. This eliminates the need for the cumbersome construction of MLWFs while strictly preserving the Hamiltonian’s symmetry, making it particularly well-suited for high-throughput workflows in electronic structure calculations of various materials. Currently, PYATB mainly offers three functional modules, i.e., Bands, Geometric, and Optical. This allows ABACUS, in combination with PYATB, to compute basic band structures, fat bands, and projected density of states (PDOS). Additionally, it can analyze the Berry curvature and Chern number of topological materialsXiao, Chang, and Niu (2010); Hasan and Kane (2010), as well as compute first- and second-order optical responses such as optical conductivity, shift current, second harmonic generation Sturman and Fridkin (1992); Sipe and Shkrebtii (2000); Jin and He and Berry curvature dipoleSodemann and Fu (2015); Pang, Jin, and He .
Once the Hamiltonian matrix $\bm{H}(\mathbf{k})$ and the overlap matrix $\bm{S}(\mathbf{k})$ are given, the element dipole matrix is given by
$A_{\mu\nu}^{\mathbf{R}}(\mathbf{k})=\sum_{\mathbf{R}}\mathrm{e}^{i\mathbf{k}\cdot\mathbf{R}}\langle\phi_{\mu 0}|\mathbf{r}|\phi_{\nu\mathbf{R}}\rangle.$ (114)
Using these input parameters, PYATB solves the following generalized eigenvalue problem of Eq. 82. Subsequently, based on the wavefunctions, the Berry curvatureXiao, Chang, and Niu (2010); Jin, Zheng, and He (2021) can be calculated. This allows for the calculation of various topological and optical properties of materials. One can refer to Ref. Jin et al., 2023 for the implementation details.
VI.7 Hefei-NAMD
@@
Hefei Non-Adiabatic Molecular Dynamics (Hefei-NAMD) is an ab initio simulation suite for studying excited carrier dynamics in condensed matter systems. It combines real-time time-dependent density functional theory (TDDFT) with the fewest-switches surface hopping scheme and classical-path approximation. Hefei-NAMD has been used to investigate processes such as charge transfer, electron–hole recombination, spin dynamics, and exciton dynamics Zheng et al. (2019); Zheng, Zheng, and Zhao (2022); Jiang et al. (2021); Zheng et al. (2023). It allows the study of excited carrier dynamics in energy, real, and momentum spaces, while also exploring interactions with phonons, defects, and molecular adsorptions, providing valuable insights into ultrafast dynamics at the atomic scale. Hefei-NAMD works by interfacing with other ab initio codes, e.g. VASP, ABACUS, etc. Figure 34 shows the flowchart of the simulation using Hefei-NAMD interfaced with ABACUS.
@@
VI.8 PEXSI
PEXSI (Pole EXpansion and Selected Inversion) Lin et al. (2009, 2013) offers an efficient approach for solving the KS equation while circumventing the computational limitations inherent in conventional diagonalization techniques. This method directly computes the real-space density matrix through a $P$-term pole expansion of the Fermi-Dirac function, expressed as
$f(\bm{H}-\mu\bm{S})\approx\mathrm{Im}\sum_{l=1}^{P}\omega_{l}^{\rho}\left[\bm{H}-(z_{l}+\mu)\bm{S}\right]^{-1},$ (115)
where $\bm{H}$ denotes the Hamiltonian matrix, $\bm{S}$ is the overlap matrix, $z_{l}$ and $\omega_{l}$ correspond to the precomputed poles and weights of the complex contour integration, respectively. The numerical stability of this expansion is ensured through careful selection of these parameters. The computationally intensive matrix inversion operations are efficiently performed using the selected inversion algorithm.
@@ VI.8 PEXSI
Notably, PEXSI exhibits superior computational scaling compared to traditional diagonalization methods. While standard diagonalization scales cubically as $O(N^{3})$, PEXSI achieves at most quadratic scaling ($O(N^{2})$), where $N$ is the dimension of the matrix The method’s architecture features a two-level parallelization scheme, incorporating both pole expansion parallelism and selected inversion parallelism, thereby demonstrating exceptional scalability. ABACUS has implemented an interface to the parallel version of PEXSI for solving the KS equation expanded with the NAO basis.
VI.9 Other Packages
ASE (Atomic Simulation Environment) comprises a comprehensive suite of tools and Python modules designed for setting up, manipulating, executing, visualizing, and analyzing atomic simulations ASE (2024a). We have developed an ABACUS calculator ASE (2024b) that interfaces with ASE (version 3.23.0b1) and that enables seamless integration of ASE’s robust functionalities for both pre-processing and post-processing tasks. Regarding pre-processing capabilities, the interface facilitates tasks such as converting structural files into various formats, generating $k$-point grids, etc. For post-processing, the integration with ASE empowers ABACUS users with advanced tools for visualizing and analyzing simulation results. Moreover, the self-consistent calculations of ABACUS can be combined with ASE’s built-in optimization algorithms, facilitating structural relaxation procedures, conducting MD simulations under varying conditions, performing precise phonon calculations based on the finite displacement method, carrying-out global optimization by genetic algorithm (GA), locating transition states (TS) of reactions by saddle point refinement, etc. For TS search, as an example, the nudged elastic band (NEB) Jónsson, Mills, and Jacobsen (1998); Henkelman and Jónsson (2000); Henkelman, Uberuaga, and Jónsson (2000) method and the Dimer Henkelman and Jónsson (1999); Olsen et al. (2004); Heyden, Bell, and Keil (2005); Kästner and Sherwood (2008) method, which represent double-ended TS searching methods and single-end TS methods respectively, are implemented in ASE. Besides, some other enhanced NEB methods are coded in the ASE package, such as dynamic NEB Lindgren, Kastlunger, and Peterson (2019) and AutoNEB Kolsbjerg, Groves, and Hammer (2016). Furthermore, there are other saddle point refinement algorithms implemented in ASE, and some of them show better performance, like Sella Hermes et al. (2019, 2021, 2022). With the above-mentioned methods available in ASE, ABACUS can be used for searching for transition states.
ATST-Tools scripts suite ATS (2024) is prepared for handy usage of ABACUS and ASE in TS locating jobs (see workflow in Fig. 35). Apart from directly using TS search tools in ASE, the cooperation of different TS search methods can be done due to ASE’s coding flexibility, which leads to better TS locating functionality. One way is to first generate a rough TS by NEB, then utilize a single-ended method like Dimer or Sella to optimize the TS in the target threshold, which incorporates the advantage of both methods.
@@ VI.9 Other Packages
Phonopy Togo et al. (2023); Togo (2023) is a versatile open-source package designed for calculating phonon and related lattice dynamics properties from first-principles calculations. With the addition of the ABACUS interface in Phonopy version 2.19.1, users can now adopt ABACUS to compute the electronic structures and force constants that Phonopy requires for its phonon calculations. Wang et al. (2020); Liu et al. (2024b)
ShengBTE Li et al. (2014) is an open-source package aimed to solve the linearized Boltzmann Transport Equation (BTE) for phonons, enabling the ab initio prediction of lattice thermal conductivity in bulk crystalline solids and nanowires. It goes beyond the relaxation-time approximation by explicitly incorporating three-phonon scattering processes and isotopic scattering, yielding predictive results without empirical fitting. ShengBTE requires inputs of second- and third-order interatomic force constants (IFCs), traditionally computed using DFT packages, often aided by preprocessing tools. ABACUS can provide inputs for ShengBTE starting from v3.2.3.
@@
TB2JHe et al. (2021) is an open-source Python package for automatic computation of magnetic interactions between atoms of magnetic crystals from density functional Hamiltonians based on Wannier functions or linear combinations of atomic orbitals. The program is based on Green’s function method with the local rigid spin rotation treated as a perturbation Liechtenstein et al. (1987). The ABACUS interface has been added since TB2J version 0.8.0. With the Hamiltonian matrix $\mathbf{H}(\mathbf{k})$ and the overlap matrix $\mathbf{S}(\mathbf{k})$ obtained by an LCAO calculation, TB2J is able to compute the isotropic exchange, symmetric anisotropic exchange, and the Dzyaloshinskii-Moriya interaction (DMI). In 2024, Zhang et al. Zhang et al. (2024e) study GeFe3N by using ABACUS-TB2J, which help them study the short-range order and strong interplay between local and itinerant magnetism in GeFe3N.
@@
Wannier90 Pizzi et al. (2020) is a software designed for calculating maximally-localized Wannier functions (MLWFs) Marzari and Vanderbilt (1997), which facilitates cost-effective band structure calculations. MLWFs are often used to construct model Hamiltonians. The post-processing capabilities of Wannier90 enable the calculation of various properties, such as Berry curvature, Berry curvature dipole, and shift currentPizzi et al. (2020). Additionally, MLWFs generated by Wannier90 can be analyzed for material surface states and topological properties using the Wannier Tools Wu et al. (2018). According to the standard scheme for generating maximally localized Wannier functions in Wannier90, ABACUS has implemented an interface compatible with the Wannier90 software. This interface allows ABACUS to generate the necessary files for Wannier90, including overlap files between the periodic parts of Bloch functions at neighboring $\mathbf{k}$-points, projection files of Bloch functions onto trial localized orbitals, eigenvalue files of the energy bands, and real-space distribution files of the periodic parts of Bloch functions.
APEX (Alloy Property EXplorer) Li et al. (2025) is versatile and extensible Python package for general alloy property calculations. This package enables users to conveniently establish a wide range of cloud-native property-test workflows by utilizing ABACUS. APEX currently supports the calculations including equation of state (EOS), elastic constants, surface energy, interstitial formation energy, vacancy formation energy, generalized stacking fault energy (Gamma line), and phonon spectra.
CALYPSO (Crystal structure AnaLYsis by Particle Swarm Optimization) Wang et al. (2012) is a software designed for prediction of stable and metastable crystal structures. It employs advanced techniques like particle swarm optimization (PSO) to efficiently explore complex potential energy surfaces. CALYPSO has an interface with ABACUS to compute accurate energies and forces during the structure search process. The combined CALYPSO-ABACUS workflow can be utilized for high-throughput structure prediction and inverse design of functional materials under specified chemical compositions and external conditions like pressure.
USPEX (Universal Structure Predictor: Evolutionary Xtallography) Oganov and Glass (2006) represents a powerful methodology for crystal structure prediction. USPEX employs evolutionary algorithms to efficiently predict stable and metastable structures solely on chemical composition. USPEX can be utilized with ABACUS for accurate energy and force calculations during evolutionary structure searches.
VII Summary